# devoted to memory of Alexei Alexandrovich Starobinsky

### v15

@@ II.9 Hubble volume entropy vs entropy of the cosmological horizon
$S_{\rm bulk}=s_{\rm vac}V_{H}=\frac{4\pi R^{3}}{3}s_{\rm vac}=\frac{\pi}{GH^{2}}=4\pi KA=S_{\rm hor}\,,$ (18)
@@
The Hubble-volume entropy coincides with the Gibbons-Hawking entropy of the cosmological horizon. However, here it is the thermodynamic entropy coming from the local entropy of the de Sitter quantum vacuum, rather than the entropy of the horizon degrees of freedom. It is typically assumed that the horizon microstates are concentrated in the region of the Planck length $l_{\rm Pl}=1/M_{\rm Pl}$ in the vicinity of horizon, and this leads to the area law for the total entropy inside the horizon.Padmanabhan2010 The local thermodynamics leads to the same area law.
@@
The connection between the bulk entropy of the region inside the horizon due to the local vacuum thermodynamics and the surface entropy of the horizon has the holographic origin, which is related to the horizon as null surface.Padmanabhan2010 This is the reason why the presence of the horizon allows to know the total entropy of the Hubble volume without knowledge on the local entropy inside the horizon.
@@
This bulk-surface correspondence in Eq.(18) is valid only in the ($3+1$)-dimension. In the general $d+1$ dimension of spacetime, the same approach gives the factor $(d-1)/2$ in the relation between the entropy of the Hubble volume and the Gibbons-Hawking entropy of the cosmological horizon, $S_{\rm bulk}=\frac{d-1}{2}\,S_{\rm hor}$. This may add to the peculiarities of the $d=3$ space dimension,Barrow1983 where in particular the mass dimension of the gravitational coupling, $[K]=d-1$, coincides with the mass dimension of curvature, $[{\cal R}]=2$. The same concerns such pair of thermodynamically conjugate variables as electric field with the mass dimension $[E]=2$ and the electric induction with mass dimension $[D]=d-1$. Their dimensions also coincide only for $d=3$. Discussion of the natural dimensions of physical quantities in $d=3$ can be found in Ref. Volovik2022f .
III Hubble volume vs the volume of Universe
III.1 Thermal fluctuations of de Sitter state
It is not excluded that our Universe is finite. Its volume $V$ might be comparatively small, not much larger than the currently observed Hubble volume $V_{H}$.Lehners2024
If the Universe is finite and if the de Sitter state represents the excited thermal state of the quantum vacuum, the thermal fluctuations of the deep quantum vacuum may become important. According to Landau-Lifshitz,LandauLifshitz the thermal fluctuations are determined by the compressibility of the system and by its volume. In case of the Universe with the volume $V$, the fluctuations of the vacuum energy density are given by:Volovik2013
Note the main difference between the thermal fluctuations and quantum fluctuations. The contribution of the quantum fluctuations of the relativistic quantum fields to the vacuum energy density is typically on the order of $M_{\rm Pl}^{4}$, where $M_{\rm Pl}$ is the Planck mass. But in the fully equilibrium vacuum state this contribution is cancelled by the ultraviolet trans-Planckian degrees of freedom due to the thermodynamic Gibbs-Duhem relation.KlinkhamerVolovik2008 ; Volovik2013 This cancellation is universal, being valid both for the relativistic vacuum states and for the non-relativistic grounds states of the condensed matter systems. The contribution of thermal fluctuations to the vacuum energy is in the range of the infrared physics, which is determined by temperature $T$.
III.2 Vacuum compressibility
The vacuum compressibility is not constrained by the thermodynamic laws and is determined by the ultraviolet physicsKlinkhamerVolovik2008 with its Planck energy scale, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$. In this relation it is similar to the gravitational coupling, which is also determined by the Planck scale, $K\sim M_{\rm Pl}^{2}$. On the other hand, the temperature corrections to $\chi_{\rm vac}^{-1}$ and $K$ as well as the Casimir corrections are within the range of the infrared physics, see Ref. VolovikZelnikov2003 for the universal temperature correction to the gravitational coupling $K$. The contribution to the compressibility from the infrared range was discussed in Refs.BLHu2022 ; BLHu2024 ; BLHu2024a ; Widom1998 . The negative contributions to compressibility obtained in these papers do not violate the stability of the quantum vacuum, since these contributions represent the corrections, which are small compared to the main value of the vacuum compressibility, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$.
In the excited vacuum – the de Sitter state with the temperature $T=H/\pi$ and the energy density $<\epsilon_{\rm vac}>\sim M_{\rm Pl}^{2}H^{2}$ – the relative magnitude of thermal fluctuations is determined by the ratio of the Hubble volume to the volume of the Universe:
[…]
The volume of the present Universe exceeds the Hubble volume, $V>V_{H}$, and thus the thermal fluctuations of the vacuum energy density are still relatively small.
@@ IV.1 Multi-metric gravity
Since the heat exchange between the bodies or between the systems is the main source of the emergent thermodynamics,Strasberg2024 we can consider the de Sitter thermodynamics from the point of view of the heat transfer between different cosmological objects.
Such heat exchange can be discussed in the frame of the so-called multi-metric gravity, see Ref.Wood2024 and references therein. The corresponding model action of the whole system can be written as the sum of actions of the sub-systems in the same coordinate spacetime:
@@
Following Frogatt and NielsenFrogattNielsen1991 one can introduce $N$ independent tetrad fields $e^{a(n)}_{\mu}$ for $N$ fermionic species. In this multi-tetrad gravity one has:
${\cal L}_{n}=e_{abcd}K_{n}\,R^{ab(n)}\wedge e^{c(n)}\wedge e^{d(n)}+$
$+e_{abcd}\Lambda_{n}\,e^{a(n)}\wedge e^{b(n)}\wedge e^{c(n)}\wedge e^{d(n)}\,,$ (22)
with the corresponding action for the fermionic species:
@@
On the other hand, gravity with multiple tetrad fields may come from the Akama-Diakonov-Wetterich theory, Akama1978 ; Wetterich2004 ; Diakonov2011 ; VladimirovDiakonov2012 ; VladimirovDiakonov2014 ; Sindoni2012 ; ObukhovHehl2012 ; Volovik1990 ; WeiLu2024 where the tetrads are formed as composite objects – the bilinear combinations of the fundamental fermionic fields:
@@ IV.2 Heat exchange in multi-metric gravity
The heat exchange between the sub-Universes leads to their equilibration with formation of the common expansion rate and thus the common temperature. We consider first the system of two Universes, assuming that in each of them the entropy of horizon obeys the area law, and show that the maximum entropy corresponds to the state in which both states have the same expansion rate.
@@
[…]
@@
$S=-\int d^{4}x\sqrt{-g_{(1)}}\left(K_{1}{\cal R}\{g_{\mu\nu(1)}\}+\Lambda_{1}\right)-$
$-\int d^{4}x\sqrt{-g_{(2)}}\left(K_{2}{\cal R}\{g_{\mu\nu(2)}\}+\Lambda_{2}\right)\,.$ (29)
@@
Let us now allow for the energy exchange (the heat exchange) between these two Universes (analogs of the two layers of graphene). This exchange can be realized by the matter field, which interacts with both metrics. It leads to the variations of the Hubble parameters $H_{1}$ and $H_{2}$ at fixed $E_{V}$. If we ignore the thermalization of matter by de Sitter environment, the heat exchange will finally produce the equilibrium state with the maximum entropy $S$, in which the Hubble parameters become equal:
@@
In case of the arbitrary number $N$ of the sub-Universes, the heat exchange between them leads to the state of the Universe in which all the sub-Universes coherently expand with the same rate $H$, i.e., with the same de Sitter metric in all the subsystems. The whole Universe has the gravitational coupling $K$ equal to the sum of the individual couplings in the sub-Universes and the vacuum energy density equal to the sum of energy densities of subsystems:
$E_{V}=6KH^{2}V\,\,,\,\,K=\sum_{n}K_{n}\,\,,\,\,\Lambda=\sum_{n}\Lambda_{n}\,,$ (34)
with $T_{n}=T=H/\pi$.
@@ IV.3 Thermodynamics from the multi-metric ensemble
For large $N$, the random distribution of the parameters results in the exponential behaviour of the distribution functions, $w_{n}\propto\exp(-E_{n}/T)$, with the same parameter $T$ for all subsystems. As in the statistical ensemble of atoms, where the temperature of the system is determined by the physical processes, the temperature of the ensemble of the sub-Universes factor is also determined by the physical processes. In our case it is the of the behaviour of matter (atom) in the de Sitter environment, which gives $T=H/\pi$. This connection between $T$ and $H$ is rather natural. Both, the parameter $T$, which plays the role of temperature, and the Hubble parameter $H$ are the quantities which in equilibrium become common for all the subsystems in the ensemble, and they have the same dimension of inverse time, $[T]=[H]=[1/t]$.
@@ IV.4 Regularization vs thermalization
The multi-metric ensemble may include the ensemble of $N$ species of Weyl or Dirac fermions. At large $N$, all tetrads in the random ensemble approach the same value, $e^{a(n)}_{\mu}\rightarrow e^{a}_{\mu}$, and thus in the equilibrium state all fermionic species experience the same geometry. In Refs. FrogattNielsen1991 ; ChadhaNielsen1983 the formation of the common Lorentz invariance for different fermionic species is also considered. But this is achieved by the renormalization group effect in the infrared limit, instead of thermalization. This suggests the possible connection between renormalization and thermalization.
One may expect that the heat exchange between subsystems leads not only to the coherence of the de Sitter states, but to the general coherence of the metric fields, when the metric fields $g_{\mu\nu(n)}$ of the subsystems become equal, thus forming the common metric $g_{\mu\nu}$. If this is true, this could be some kind of the thermodynamic gravity, but without using the holographic principle.
@@ IV.5 Coherence vs thermalization
At first glance this formation of the time dependent coherent state looks similar to the formation of the coherent spin precession (magnon BEC) from the incoherent precessions about the local magnetic fields with random frequencies $\omega_{n}$.BunkovVolovik2013 The latter corresponds to random $H_{n}$. The coherence develops due to the spin currents between the regions of local precessions – the analog of the heat exchange. The common frequency $\omega$ of the coherent precession corresponds to the common Hubble parameter $H$. They have the same dimension, $[H]=[\omega]=[1/t]$, while the dimensionless magnon number (or the spin projection $S_{z}$) corresponds to the entropy which is also dimensionless. However, the source of the coherence is the exchange of spin instead of exchange of the energies, and also the formation of coherent state is due to minimization of the total energy $E$ at fixed projection $S_{z}$ of the total spin on magnetic field. This leads to the common frequency, which plays the role of the chemical potential $\mu$, which emerges due to the exchange of particles between the regions. That is why this process is quite opposite to the formation of temperature and entropy, where energy is fixed, while entropy reaches maximum value. This process corresponds to the minimization of energy at fixed total entropy.
Anyway, in all the cases the parameters, which are the same in all subsystems in equilibrium – common frequency $\omega$, common chemical potential $\mu$, common temperature $T$, common angular velocity $\boldsymbol{\Omega}$ and common Hubble parameter $H$ – have the same dimension, $[T]=[H]=[\mu]=[\omega]=[\Omega]=[1/t]$. Note also that together with the variables $T$, $\mu$ and $\boldsymbol{\Omega}$, the Hubble parameter $H$ enters the chiral anomaly effects, see Sec. IV.6.
@@
Let us consider the chiral vortical effect – the appearance of chiral current in fermionic system in the presence of rotation. For Dirac fermions in the flat spacetime the chiral current ${\bf j}_{A}$ contains the following contributions:Vilenkin1979 ; Vilenkin1980 ; Kharzeev2016 ; Stone2018 ; Abramchuk2024
@@
In the de Sitter state the curvature term can be added.Flachi2018 ; Khakimov2024 According to Ref. Khakimov2024 the gradient expansion gives:
@@
The same term is obtained for the anti-de Sitter spacetime where $H^{2}<0$.Ambrus2021 In Ref.Flachi2018 the curvature term comes from the shift of the fermionic mass gap due to curvature, and this leads to the Eq.(36) with opposite sign.

### v16

@@ III.4 Hubble volume entropy vs entropy of the cosmological horizon
$S_{\rm bulk}=s_{\rm vac}V_{H}=\frac{4\pi R^{3}}{3}s_{\rm vac}=\frac{\pi}{GH^{2}}=\frac{A}{4G}=S_{\rm hor}\,,$ (27)
@@
The Hubble-volume entropy coincides with the entropy attributed to the cosmological horizon as suggested by Gibbons and Hawking. However, Eq.(27) demonstrates that the thermodynamic entropy comes from the local entropy of the de Sitter quantum vacuum, rather than from the hypothetical horizon degrees of freedom. In this hypothesis it assumed that there are the horizon microstates, which are concentrated in the region of the Planck length $l_{\rm Pl}=1/M_{\rm Pl}$ in the vicinity of horizon, and this leads to the area law for the total entropy inside the horizon, see e.g. Ref. Padmanabhan2010 . The local thermodynamics in the de Sitter state leads to the same area law, but without assumptions on the spatial distribution of the degrees of freedom.
@@
The connection between the bulk entropy of the region inside the horizon due to the local vacuum thermodynamics and the surface entropy of the horizon may have the holographic origin, when the horizon is considered as null surface.Padmanabhan2010 This may be the reason why the presence of the horizon allows us to know the total entropy of the Hubble volume without knowledge on the distribution of the local entropy inside the horizon.
@@
However, it appears that the bulk-surface correspondence in Eq.(27) is valid only in the ($3+1$)-dimensional spacetime. In the general $d+1$ dimension of spacetime, the same approach gives the factor $(d-1)/2$ in the relation between the entropy of the Hubble volume and the Gibbons-Hawking entropy of the cosmological horizon, $S_{\rm bulk}=\frac{d-1}{2}\,S_{\rm hor}$. This may add to the peculiarities of the $d=3$ space dimension,Barrow1983 where in particular the mass dimension of the gravitational coupling, $[K]=d-1$, coincides with the mass dimension of curvature, $[{\cal R}]=2$. The same concerns such pair of thermodynamically conjugate variables as electric field with the mass dimension $[E]=2$ and the electric induction with mass dimension $[D]=d-1$. Their dimensions also coincide only for $d=3$. Discussion of the natural dimensions of physical quantities in $d=3$ can be found in Ref. Volovik2022f .
@@ IV.1 Multi-metric gravity
As we know, the main source of the emergent thermodynamics is the heat exchange between the bodies or between the systems, see also Ref.Strasberg2024 . Following this rule, we can consider the de Sitter thermodynamics from the point of view of the heat transfer between different cosmological objects or different Universes.
The heat exchange can be discussed in the frame of the so-called multi-metric gravity, see Ref.Wood2024 and references therein. The corresponding model action of the whole system can be written as the sum of actions of the sub-systems in the same coordinate spacetime:
@@
Following Froggatt and NielsenFroggattNielsen1991 one can introduce $N$ independent tetrad fields $e^{a(n)}_{\mu}$ for $N$ fermionic species. In this multi-tetrad gravity one has the ensemble of the gravitational actions:
${\cal L}_{n}=e_{abcd}\left(K_{n}\,R^{ab(n)}\wedge e^{c(n)}\wedge e^{d(n)}+\Lambda_{n}\,e^{a(n)}\wedge e^{b(n)}\wedge e^{c(n)}\wedge e^{d(n)}\right)\,,$ (29)
and the corresponding ensemble of actions for the fermionic species:
@@
Gravity with multiple tetrad fields may also come from the Akama-Diakonov-Wetterich theory, Akama1978 ; Wetterich2004 ; Diakonov2011 ; VladimirovDiakonov2012 ; VladimirovDiakonov2014 ; Sindoni2012 ; ObukhovHehl2012 ; Volovik1990 ; WeiLu2024 where the tetrads are formed as composite objects – the bilinear combinations of the fundamental fermionic fields:
@@
On the levels of particles, the vestigial gravity acts with different strength on fermions and bosons. In principle, it is not excluded that such gravity can be formed at an early stage of the development of the Universe.
@@ IV.2 Heat exchange in multi-metric gravity
The heat exchange between the sub-Universes leads to their equilibration with formation of the common expansion rate and thus the common temperature. We consider first the system of two sub-Universes, assuming that in each of them the entropy of horizon obeys the area law, and show that the maximum entropy corresponds to the situation in which both states acquire the same expansion rate.
@@
Let us assume that the bulk-surface correspondence is valid, i.e. the entropy of the Hubble volume $V_{H}$ is equal to the Gibbons-Hawking entropy of cosmological horizon, $S_{V_{H}}=S_{\rm hor}=4\pi KA$. Then the total entropy $S_{V}$ in the volume $V\gg V_{H}$ can be obtained from the entropy of the Hubble volume $V_{H}$:
@@
$S=-\int d^{4}x\sqrt{-g_{(1)}}\left(K_{1}{\cal R}\{g_{\mu\nu(1)}\}+\Lambda_{1}\right)-\int d^{4}x\sqrt{-g_{(2)}}\left(K_{2}{\cal R}\{g_{\mu\nu(2)}\}+\Lambda_{2}\right)\,.$ (36)
@@
Let us now allow for the energy exchange (the heat exchange) between these two sub-Universes (analogs of the two layers of graphene). This exchange can be realized by the matter field, which interacts with both metrics. It leads to the variations of the Hubble parameters $H_{1}$ and $H_{2}$ at fixed $E_{V}$. If we ignore the thermalization of matter by de Sitter environment, the heat exchange will finally produce the equilibrium state with the maximum entropy $S$, in which the Hubble parameters become equal:
@@
In case of the arbitrary number $N$ of the sub-Universes, the heat exchange between them leads to the equilibrium state of the Universe in which all the sub-Universes coherently expand with the same rate $H$, i.e., with the same de Sitter metric in all the subsystems. In this equilibrium Universe, the gravitational coupling $K$ is equal to the sum of the individual couplings in the sub-Universes and the vacuum energy density is equal to the sum of energy densities of subsystems:
$E_{V}=6KH^{2}V\,\,,\,\,K=\sum_{n}K_{n}\,\,,\,\,\Lambda=\sum_{n}\Lambda_{n}\,.$ (41)
All the substates acquire the same temperature in equilibrium, $T_{n}=T=H/\pi$.
@@ IV.3 Thermodynamics from the multi-metric ensemble
For large $N$, the random distribution of the parameters results in the exponential behaviour of the distribution functions, $w_{n}\propto\exp(-E_{n}/T)$, with the same parameter $T$ for all subsystems. As in the statistical ensemble of atoms, where the temperature of the system is determined by the physical processes, the temperature of the ensemble of the sub-Universes is also determined by the physical processes. In our case it is the behaviour of matter (atom) in the de Sitter environment, which gives $T=H/\pi$. This connection between $T$ and $H$ is rather natural. Both, the parameter $T$, which plays the role of temperature, and the Hubble parameter $H$ are the quantities which in equilibrium become common for all the subsystems in the ensemble, and they have the same dimension of inverse time, $[T]=[H]=[1/t]$.
@@ IV.4 Regularization vs thermalization
The multi-metric ensemble may include the ensemble of $N$ species of Weyl or Dirac fermions. At large $N$, all tetrads in the random ensemble approach the same value, $e^{a(n)}_{\mu}\rightarrow e^{a}_{\mu}$, and thus in the equilibrium state all fermionic species experience the same geometry. In Refs. FroggattNielsen1991 ; ChadhaNielsen1983 the formation of the common Lorentz invariance for different fermionic species was also considered. But this was achieved by the renormalization group effect in the infrared limit, instead of thermalization. This suggests the possible connection between renormalization and thermalization.
One may expect that the heat exchange between subsystems leads not only to the coherence of the de Sitter states, but to the general coherence of the metric fields, when the metric fields $g_{\mu\nu(n)}$ of the subsystems become equal, thus forming the common metric $g_{\mu\nu}$. If this is true, this could be a kind of the thermodynamic gravity, but without using the holographic principle.
@@ IV.5 Coherence vs thermalization
At first glance this formation of the coherent de Sitter expansion from the ensemble of the random microstates looks similar to the formation of the Bose-Einstein condensate (BEC) of magnons.BunkovVolovik2013 The magnon BEC represents the coherent precession of all spins, which results from the incoherent precessions of the individual spins with random frequencies $\omega_{n}$ in the local magnetic fields. The originally random frequencies correspond to the random Hubble parameters $H_{n}$ in different sub-Universes. The coherence of precession develops due to the spin currents between the regions of the local precessions – the analog of the heat exchange between the sub-Universes. The formed common frequency $\omega$ of the coherent precession corresponds to the formed common Hubble parameter $H$ of the large Universe. Moreover, they have the same dimension of the inverse time, $[H]=[\omega]=[1/t]$, while the dimensionless number of magnons (or, which is the same, the dimensionless spin projection $S_{z}$) corresponds to the total entropy of de Sitter which is also dimensionless.
However, the source of the coherence is the exchange of spins instead of the exchange of the energies. As a result the formation of the coherent state of spin precession is due to minimization of the total energy $E$ at fixed projection $S_{z}$ of the total spin on magnetic field. This leads to the common frequency of precession. In the magnon BEC interpretation, the common frequency plays the role of the chemical potential $\mu$ for magnons, which becomes constant in space due to the exchange of magnons between the regions. That is why this process is quite opposite to the formation of the common temperature, where the total energy is fixed, while the total entropy reaches its maximum value.
Anyway, in all the cases the parameters, which are the same in all subsystems in equilibrium – common frequency $\omega$, common chemical potential $\mu$, common temperature $T$, common angular velocity $\boldsymbol{\Omega}$ and now also the common Hubble parameter $H$ – all of them have the same dimension of inverse time, $[T]=[H]=[\mu]=[\omega]=[\Omega]=[1/t]$. In Sec. IV.6 it will be shown that the Hubble parameter $H$ enters the chiral anomaly effects together with the other thermodynamic variables: $T$, $\mu$ and $\boldsymbol{\Omega}$. All this supports the thermodynamic nature of the Hubble parameter.
@@
As the thermodynamic quantity, the Hubble parameter $H$ participates in different thermodynamical effects. We consider this using as an example the chiral vortical effect – the appearance of the chiral current in fermionic systems in the presence of rotation. For Dirac fermions in the flat spacetime, the chiral current ${\bf j}_{A}$ contains the following contributions from temperature, chemical potential and angular velocity:Vilenkin1979 ; Vilenkin1980 ; Kharzeev2016 ; Stone2018 ; Abramchuk2024
@@
In the de Sitter state the gravitational thermodynamical variable – the curvature ${\cal R}$ – is added.Flachi2018 ; Khakimov2024 According to Ref. Khakimov2024 the gradient expansion gives the following contribution to the chiral current, which is very similar to the contributions of other thermodynamic quantities in Eq.(42):
@@
The same term is obtained for the anti-de Sitter spacetime where $H^{2}<0$.Ambrus2021 On the other hand, in Ref.Flachi2018 the curvature term comes from the shift of the fermionic mass gap, $-M^{2}\rightarrow-(M^{2}+H^{2})$, and this leads to the Eq.(43) with opposite sign.